A rock has a mass of 0.3kg. While being swung in a sling the rock has an angular momentum of 8.1kg*m2/s. When the rock is released, it has a height of 2.2m. At this moment the rock it traveling horizontally. The rock lands 24m from where it was launched. How long was the sling? Express your answer in meters.
1
Expert's answer
2020-04-30T10:41:18-0400

Mass of rock = "0.3 kg"

Angular momentum = "8.1 kgm2\/s"

Height = "2.2 m"

Distance = "24 m"

"\because" Angular momentum = "iw"

"w= v\/r"

"v^2= u^2+2as"

"v^2= 0^2+2(9.8\u00d72.2)"

"v^2= 43.12"

"v= 6.566 ml"

Therefore "w= 6.566\/24 =0.2736 rad\/s"

"8.1=i (0.2736)"

"i=8.1\/0.2736 =29.605 m"

The sling was "29.605 m"